OpenStudy (ttop0816):

Please help! Solve for x for 180° < x < 270°. sin^2(x) = 1/4

OpenStudy (linknissan):

Ok Let me help! 180-270.. Andle is in the third quadrant.. Right? Waht is the sign of X and Y quadrant in 3rd Quarant?

OpenStudy (linknissan):

Angle*

OpenStudy (lncognlto):

sin^2(x) = 1/4 therefore sin (x) = square root (1/4) Therefore sin (x) = (-1/2) or (1/2)

OpenStudy (ttop0816):

so what would the degree be??